# FactAppeal: Identifying Epistemic Factual Appeals in News Media

### v1

@@ Abstract
How is a factual claim made credible? We propose the novel task of Epistemic Appeal Identification, which identifies whether and how factual statements have been anchored by external sources or evidence. To advance research on this task, we present FactAppeal, a manually annotated dataset of 3,226 English-language news sentences. Unlike prior resources that focus solely on claim detection and verification, FactAppeal identifies the nuanced epistemic structures and evidentiary basis underlying these claims and used to support them. FactAppeal contains span-level annotations which identify factual statements and mentions of sources on which they rely. Moreover, the annotations include fine-grained characteristics of factual appeals such as the type of source (e.g. Active Participant, Witness, Expert, Direct Evidence), whether it is mentioned by name, mentions of the source’s role and epistemic credentials, attribution to the source via direct or indirect quotation, and other features. We model the task with a range of encoder models and generative decoder models in the 2B-9B parameter range. Our best performing model, based on Gemma 2 9B, achieves a macro-$F_{1}$ score of 0.73.111Data available on Github, cc-by-4.0 license.
@@
Guy Mor-Lan, Tamir Sheafer and Shaul R. Shenhav Hebrew University of Jerusalem {guy.mor, tamir.sheafer, shaul.shenhav}@mail.huji.ac.il
@@ 1 Introduction
In an era marked by pervasive misinformation and heightened skepticism of media reporting, understanding how factual claims are presented has become more important than ever.While research in claim detection and verification has made substantial progress Sauri and Pustejovsky (2009); Thorne et al. (2018); Hassan et al. (2017); Wadden et al. (2020); Aly et al. (2021), most existing methods focus on the content of the statements in isolation and overlook the epistemic structures that confer credibility and persuasive force to these claims. In news media, for example, the credibility of a claim is not only determined by its content but also by the way it appeals to external sources of knowledge—be it through expert testimony, official statements, or direct empirical evidence. Understanding how factual claims are anchored by appeals to external sources is also important for broader tasks in discourse analysis, fact-checking, and the study of knowledge flows in the media.
@@ 1.1 Epistemic Appeal Identification
We propose the task of Epistemic Appeal Identification, which requires determining whether a sentence presents a factual claim and, if so, identifying how it invokes an external source or evidence to support that claim. This task requires identifying the source of epistemic authority, as well as classifying of the type and method of appeal. This new task pushes the boundaries of traditional factuality detection by introducing a rich layer of epistemic reasoning, crucial for understanding how information is conveyed and validated in public discourse.
@@ 2.1 Verifiability and Claim Verification
Early work on factuality detection aimed at determining whether statements describe verifiable events Sauri and Pustejovsky (2009, 2012); Hassan et al. (2017). More recent lines of research have emphasized claim verification, exemplified by large-scale benchmarks such as FEVER Thorne et al. (2018), which require systems to determine if a claim is supported or refuted based on evidence. Other datasets have focused on specific domains or subtasks, such as SciFact Wadden et al. (2020) for verifying scientific claims or FactRel Mor-Lan and Levi (2024) for factual entailment in news. While these resources have substantially advanced fact-checking methods, they focus primarily on detecting claims and modeling relations between claims, rather than providing a complete epistemic schema describing how a claim itself is constructed and supported.
@@ 2.2 Epistemic Modality and Argumentation
Research on epistemic modality seeks to capture linguistic markers of certainty and belief Rubin (2010); Sauri and Pustejovsky (2012), while argumentation mining explores how claims are constructed and supported within discourse Feng and Hirst (2011). A related task is epistemic stance detection, which models a source’s degree of commitment—such as certainty or doubt—toward a claim Gupta et al. (2022). Such analyses are crucial in persuasive language, where models have been trained to identify broad strategies like appeals to authority Da San Martino et al. (2019). While these works focus on belief, commitment, or general persuasive tactics, our approach offers a more granular view. By contrast, we pinpoint the concrete sources invoked (e.g., a named expert, a witness) and classify the structural nature of the appeal itself, such as the source type and method of quotation. This allows for more precise modeling of how claims receive or signal credibility.
@@ 2.3 Source Attribution and Quotation Analysis
Prior studies have addressed the task of detecting quotations and attributing them to entities Pareti (2015); Cohen et al. (2010), which is crucial for scientific, journalistic and legal texts. However, these methods do not typically classify sources by type (e.g., expert vs. witness) or capture whether appeals are invoked through direct speech or paraphrasing. We build on these works by jointly modeling factuality and source-based epistemic appeals, thereby revealing how news articles invoke or display a source’s authority to support a factual claim.
@@ 3.2 Factual Claims
We first examine the factuality of a sentence. Factual claims are sentences that primarily make a statement about the state of the external world, which could be either true or false. They correspond to what Jakobson describes as the referential function of language, which is concerned with conveying information about the external world and is “oriented toward the context” (Roman, 1960), as well as to the assertive speech act described by Searle, in which the speaker commits to the truth of what is asserted (Searle, 2013). Thus, statements that primarily convey a personal experience or subjective feeling are non-factual, and receive a null annotation:
@@
Factual appeals are factual claims accompanied by a reference to a purported source of knowledge. Appeals are generally performed via some form of reference or citation, [^2] Including unattributed quotes, in which the existence of a source is implied by its identity is not determined. which could take the form of direct quotation reproducing speech verbatim, or indirect reference including any forms of paraphrasing or knowledge mediation.
@@ 3.3 Types of Epistemic Appeals
Epistemic Appeal detected Internal or External? Human? Direct Evidence Active role? Witness Non-epistemic authority? Official Active Participant Human? Journalistic? Expert Document News Report Expert InternalExternalNoYesNoYesYesNoNoNoYesYes
@@
Officials are active participants which also have extra non-epistemic authority on events or on facts — e.g., legal, political, bureaucratic authority. Officials, such as government authorities, often provide statements that carry legal or formal weight. Importantly, officials wield power that can alter states of affairs related to the factual claim.
@@
The distinction between internal and external sources also reflects two modes or logics of epistemology —a common-wisdom logic preferring those with direct relations to the matter at hand, as opposed to an expertise-based logic preferring “detached" experts. Whereas internal sources have epistemic credentials in virtue of their specific history and contact with the situation at hand, external sources possess epistemic credo due to their attained expertise (Pierson, 1994; Collins and Evans, 2002).
@@ 4.1 Inter-Annotator Agreement Analysis
We conducted an inter-annotator agreement (IAA) analysis on a subset of data annotated by both annotators. To facilitate the comparison, each span annotation was converted into binary word-level labels. Using these labels, we computed several metrics—namely the union and intersection counts, the number of words where neither annotator marked the tag, the Intersection over Union (IoU), and Cohen’s Kappa. Table 4 summarizes the IAA statistics for each tag. The overall IoU of 0.74 and a Cohen’s Kappa of 0.82 indicate substantial agreement between the annotators. However, some span annotations are relatively rare and have few instances.
@@ 4.2 Descriptive Statistics
In Figure 6, we present the distribution of span annotations. We first observe that statements without epistemic appeals appear nearly twice as frequently as those containing appeals. We observe, moreover, that most factual appeals utilize paraphrasing (66%) rather than direct quotation (34%).
@@ 5 Experiments
Since different tag types may overlap, in this setting we represent the tags as token-level multi-label binary annotations, with 18 labels corresponding to each of the tags and possible modifier values. We fine-tune pre-trained transformer encoder models, using the base model versions of RoBERTa (Liu, 2019), DeBERTa v3 (He et al., 2021) and ModernBERT (Warner et al., 2024). The encoder models are trained for up to 12 epochs with focal loss (Lin et al., 2018).
@@
In this setting, annotations are represented as XML-style tags (similar to the presentation in Section 3). Models are trained to produce the annotated sentence given the raw sentence. We fine-tune several smaller pre-trained LLMs such as Gemma 2 (2B and 9B) (Team et al., 2024), Llama 3.1 8B (Dubey et al., 2024) and Mistral v0.3 7B (Jiang et al., 2023). The models are trained with QLORA (Dettmers et al., 2023) with 4-bit quantization for 3 epochs, with $r=256$ and $alpha=256$. We mask the loss of the input prompt and train on completions.
@@ 6 Conclusion
In this work, we introduced FactAppeal, a novel dataset and task formulation aimed at identifying epistemic appeals in news media factual claims. Our dataset captures both the factuality of claims and the underlying epistemic structures that lend these claims credibility. The experiments comparing token-level predictions using encoder models with generative LLMs underscore the challenges of modeling nuanced epistemic appeals, as well as the strength of generative models and the feasibility of sequence-to-sequence representations for this task.
@@ Appendix A Annotators
The dataset has been annotated by two annotators, one of the authors and a student research assistant receiving adequate hourly compensation. The annotators are a man and a woman in their 20s-30s from the EMEA region.

### v2

@@ Abstract
How is a factual claim made credible? We propose the novel task of Epistemic Appeal Identification, which identifies whether and how factual statements have been anchored by external sources or evidence. To advance research on this task, we present FactAppeal, a manually annotated dataset of 3,226 English-language news sentences. Unlike prior resources that focus solely on claim detection and verification, FactAppeal identifies the nuanced epistemic structures and evidentiary basis underlying these claims and used to support them. FactAppeal contains span-level annotations which identify factual statements and mentions of sources on which they rely. Moreover, the annotations include fine-grained characteristics of factual appeals such as the type of source (e.g. Active Participant, Witness, Expert, Direct Evidence), whether it is mentioned by name, mentions of the source’s role and epistemic credentials, attribution to the source via direct or indirect quotation, and other features. We model the task with a range of encoder models and generative decoder models in the 2B–9B parameter range. Our best performing model, based on Gemma 2 9B, achieves a macro-$F_{1}$ score of 0.73.111Data available at github.com/guymorlan/factappeal/, cc-by-4.0 license.
@@
Guy Mor-Lan    Tamir Sheafer    Shaul R. Shenhav Hebrew University of Jerusalem {guy.mor, tamir.sheafer, shaul.shenhav}@mail.huji.ac.il
@@ 1 Introduction
In an era marked by pervasive misinformation and heightened skepticism of media reporting, understanding how factual claims are presented has become more important than ever. While research in claim detection and verification has made substantial progress Saurí and Pustejovsky (2009); Thorne et al. (2018); Hassan et al. (2017); Wadden et al. (2020); Aly et al. (2021), most existing methods focus on the content of the statements in isolation and overlook the epistemic structures that confer credibility and persuasive force to these claims. In news media, for example, the credibility of a claim is not only determined by its content but also by the way it appeals to external sources of knowledge—be it through expert testimony, official statements, or direct empirical evidence. Understanding how factual claims are anchored by appeals to external sources is also important for broader tasks in discourse analysis, fact-checking, and the study of knowledge flows in the media.
@@ 1.1 Epistemic Appeal Identification
We propose the task of Epistemic Appeal Identification, which requires determining whether a sentence presents a factual claim and, if so, identifying how it invokes an external source or evidence to support that claim. This task requires identifying the source of epistemic authority, as well as classifying the type and method of appeal. This new task pushes the boundaries of traditional factuality detection by introducing a rich layer of epistemic reasoning, crucial for understanding how information is conveyed and validated in public discourse.
@@ 2.1 Verifiability and Claim Verification
Early work on factuality detection aimed at determining whether statements describe verifiable events Saurí and Pustejovsky (2009, 2012); Hassan et al. (2017). More recent lines of research have emphasized claim verification, exemplified by large-scale benchmarks such as FEVER Thorne et al. (2018), which require systems to determine if a claim is supported or refuted based on evidence. Other datasets have focused on specific domains or subtasks, such as SciFact Wadden et al. (2020) for verifying scientific claims or FactRel Mor-Lan and Levi (2024) for factual entailment in news. While these resources have substantially advanced fact-checking methods, they focus primarily on detecting claims and modeling relations between claims, rather than providing a complete epistemic schema describing how a claim itself is constructed and supported.
@@ 2.2 Epistemic Modality and Argumentation
Research on epistemic modality seeks to capture linguistic markers of certainty and belief Rubin (2010); Saurí and Pustejovsky (2012), while argumentation mining explores how claims are constructed and supported within discourse Feng and Hirst (2011). A related task is epistemic stance detection, which models a source’s degree of commitment—such as certainty or doubt—toward a claim Gupta et al. (2022). Such analyses are crucial in persuasive language, where models have been trained to identify broad strategies like appeals to authority Da San Martino et al. (2019). While these works focus on belief, commitment, or general persuasive tactics, our approach offers a more granular view. By contrast, we pinpoint the concrete sources invoked (e.g., a named expert, a witness) and classify the structural nature of the appeal itself, such as the source type and method of quotation. This allows for more precise modeling of how claims receive or signal credibility.
@@ 2.3 Source Attribution and Quotation Analysis
Prior studies have addressed the task of detecting quotations and attributing them to entities Pareti et al. (2013); O’Keefe et al. (2012), which is crucial for scientific, journalistic and legal texts. However, these methods do not typically classify sources by type (e.g., expert vs. witness) or capture whether appeals are invoked through direct speech or paraphrasing. We build on these works by jointly modeling factuality and source-based epistemic appeals, thereby revealing how news articles invoke or display a source’s authority to support a factual claim.
@@ 3.2 Factual Claims
We first examine the factuality of a sentence. Factual claims are sentences that primarily make a statement about the state of the external world, which could be either true or false. They correspond to what Jakobson describes as the referential function of language, which is concerned with conveying information about the external world and is “oriented toward the context” (Jakobson, 1960), as well as to the assertive speech act described by Searle, in which the speaker commits to the truth of what is asserted (Searle, 2013). Thus, statements that primarily convey a personal experience or subjective feeling are non-factual, and receive a null annotation:
@@
Factual appeals are factual claims accompanied by a reference to a purported source of knowledge. Appeals are generally performed via some form of reference or citation, [^2] Including unattributed quotes, in which the existence of a source is implied but its identity is not determined. which could take the form of direct quotation reproducing speech verbatim, or indirect reference including any forms of paraphrasing or knowledge mediation.
@@ 3.3 Types of Epistemic Appeals
EPISTEMIC APPEAL DETECTED What is the source’s relation to the event? Is the source a person? Direct Evidence Active involvement? Witness Has official authority? Active Participant Official Is the source a person? Expert Has journalistic authority? Expert Document News Report INTERNAL(Direct connection)EXTERNAL(General expertise)NOYESNOYESNOYESYESNONOYES
@@
Officials are active participants who also have extra non-epistemic authority on events or on facts — e.g., legal, political, bureaucratic authority. Officials, such as government authorities, often provide statements that carry legal or formal weight. Importantly, officials wield power that can alter states of affairs related to the factual claim.
@@
The distinction between internal and external sources also reflects two modes or logics of epistemology —a common-wisdom logic preferring those with direct relations to the matter at hand, as opposed to an expertise-based logic preferring “detached" experts. Whereas internal sources have epistemic credentials in virtue of their specific history and contact with the situation at hand, external sources possess epistemic credibility due to their attained expertise (Pierson, 1994; Collins and Evans, 2002).
@@ 4.1 Inter-Annotator Agreement Analysis
We conducted an inter-annotator agreement (IAA) analysis on a subset of data labeled by both annotators. To facilitate the comparison, each span annotation was converted into binary word-level labels. Using these labels, we computed several metrics—namely the union and intersection counts, the number of words where neither annotator marked the tag, the Intersection over Union (IoU), and Cohen’s Kappa. Table 4 summarizes the IAA statistics for each tag. The overall IoU of 0.74 and a Cohen’s Kappa of 0.82 indicate substantial agreement between the annotators. However, some span annotations are relatively rare and have few instances.
@@ 4.2 Descriptive Statistics
In Figure 6, we present the distribution of span annotations. We first observe that statements without epistemic appeals appear nearly twice as frequently as those containing appeals. Moreover, we observe that most factual appeals utilize paraphrasing (66%) rather than direct quotation (34%).
@@ 5 Experiments
Since different tag types may overlap, in this setting we represent the tags as token-level multi-label binary annotations, with 18 labels corresponding to each of the tags and possible modifier values. We fine-tune pre-trained transformer encoder models, using the base model versions of RoBERTa (Liu et al., 2019), DeBERTa v3 (He et al., 2021) and ModernBERT (Warner et al., 2024). The encoder models are trained for up to 12 epochs with focal loss (Lin et al., 2017).
@@
In this setting, annotations are represented as XML-style tags (similar to the presentation in Section 3). Models are trained to produce the annotated sentence given the raw sentence. We fine-tune several smaller pre-trained LLMs such as Gemma 2 (2B and 9B) (Gemma Team et al., 2024), Llama 3.1 8B (Grattafiori et al., 2024) and Mistral v0.3 7B (Jiang et al., 2023). The models are trained with QLORA (Dettmers et al., 2023) with 4-bit quantization for 3 epochs, with $r=256$ and $\alpha=256$. We mask the loss of the input prompt and train on completions.
@@ 6 Conclusion
In this work, we introduce FactAppeal, a novel dataset and task formulation aimed at identifying epistemic appeals in news media factual claims. Our dataset captures both the factuality of claims and the underlying epistemic structures that lend these claims credibility. The experiments comparing token-level predictions using encoder models with generative LLMs underscore the challenges of modeling nuanced epistemic appeals, as well as the strength of generative models and the feasibility of sequence-to-sequence representations for this task.
@@
Acknowledgments
This work was supported by the Israel Science Foundation (Grant no. 2501/22). We thank Hagar Kaminer for diligent research assistance, the members of the DeepStory lab, and especially Effi Levi for insightful comments and suggestions that improved the final version of this manuscript.
@@ Appendix A Annotators
The dataset has been annotated by two annotators, one of the authors and a paid student research assistant. The annotators are a man and a woman in their 20s–30s from the EMEA region.